Butch starts walking from point A at 2 miles per hour. One-half hour later, Dick starts walking from point A at 3.5 miles per hour and follows the same route. How long will it take Dick to catch up with Butch?
Answer by mananth(16949)

Butch starts walking from point A at 2 miles per hour. One-half hour later, Dick starts walking from point A at 3.5 miles per hour and follows the same route. How long will it take Dick to catch up with Butch?
..
speed of Butch = 2 mph
speed of Dick = 3.5 mph.
Dick gains 3.5 -2 = 1.5 mph.
..
Butch starts 1 1/2 hour early.
so he has traveled 3/2 * 2 = 3 miles when Dick starts.
..
dick gains 1.5 mph
..
time = distance / speed
3/1.5 = 2 hours to catch up
...
m.ananth@hotmail.ca